Please help me find the equation of the line 7/3(1/4-2/5) on line .Find the equation of the line
Found 2 solutions by stanbon, MathLover1:
Answer by stanbon(75887)

find the equation of the line
slope = 7/3
passing thru point (1/4,-2/5)
----
Form: y = mx + b
Solve for "b":
-2/5 = (7/3)(1/4) + b
----
-2/5 = 7/12 + b
-----
Mutiply thru by 60 to get:
-24 = 35 + 60b
60b = -59
b = -59/60
-----
Equation:
y = (7/3)x -(59/60)
=========================
Cheers,
Stan H.
===========

Answer by MathLover1(20855)


7/3(1/4-2/5)...it is unclear what is given...resubmit your question and state clear what is given